OpenStudy (anonymous):

Solve the equation: 4x^2 = 36

OpenStudy (anonymous):

divide both sides by 4

OpenStudy (maheshmeghwal9):

\[x^2=36/4=>x=6/2=3\]

OpenStudy (anonymous):

out of these... {6, -6} {3, -3} {2/3, -2/3} {3/2, -3/2}

OpenStudy (anonymous):

actually it will be \[\pm3\]

OpenStudy (maheshmeghwal9):

ya u asked right i did wrong

OpenStudy (maheshmeghwal9):

\[\pm3 \] is correct.

OpenStudy (anonymous):

whats that mark? before the three???

OpenStudy (maheshmeghwal9):

plus or minus mark!

OpenStudy (maheshmeghwal9):

it means \[(3,-3)\]

Parth (parthkohli):

\(\Large \color{purple}{x^{2} = 9 }\) \(\Large \color{purple}{x = \sqrt{9} }\) \(\Large \color{purple}{x = 3 }\)